# AdaZeta: Adaptive Zeroth-Order Tensor-Train Adaption for Memory-Efficient Large Language Models Fine-Tuning

### v2

@@ Abstract
Fine-tuning large language models (LLMs) has achieved remarkable performance across various natural language processing tasks, yet it demands more and more memory as model sizes keep growing. To address this issue, the recently proposed Memory-efficient Zeroth-order (MeZO) methods attempt to fine-tune LLMs using only forward passes, thereby avoiding the need for a backpropagation graph. However, significant performance drops and a high risk of divergence have limited their widespread adoption. In this paper, we propose the Adaptive Zeroth-order Tensor-Train Adaption (AdaZeta) framework, specifically designed to improve the performance and convergence of the ZO methods. To enhance dimension-dependent ZO estimation accuracy, we introduce a fast-forward, low-parameter tensorized adapter. To tackle the frequently observed divergence issue in large-scale ZO fine-tuning tasks, we propose an adaptive query number schedule that guarantees convergence. Detailed theoretical analysis and extensive experimental results on Roberta-Large and Llama-2-7B models substantiate the efficacy of our AdaZeta framework in terms of accuracy, memory efficiency, and convergence speed. [^1] Code available on GitHub https://github.com/yifanycc/AdaZeta.
@@
Yifan Yang1, Kai Zhen2, Ershad Banijamal2, Athanasios Mouchtaris2 and Zheng Zhang1 1University of California, Santa Barbara 2Amazon, Alexa AI yifanyang@cs.ucsb.edu, {kaizhen,ebanijam, mouchta}@amazon.com, zhengzhang@ece.ucsb.edu
@@ 1 Introduction
Fine-tuning large language models (LLMs) has demonstrated outstanding performance in addressing numerous natural language processing applications, such as natural language understanding Kenton and Toutanova (2019); Zhang et al. (2023), question-answering Xu et al. ; Cheng et al. (2023b), and summarization Zhang et al. (2024a). However, as the size of LLMs increases, the training process consumes progressively more GPU memory. In recent years, approaches such as quantization Dettmers et al. (2024) and parameter-efficient fine-tuning (PEFT) Hu et al. (2021); Chen et al. (2024) have been proposed to reduce memory costs during training by storing data with lower bit-depth or updating only a portion of the parameters. Despite these strategies effectively reducing memory costs, overall memory usage remains high due to the continuous reliance on a backpropagation graph.
@@
  • A new contraction method is introduced to accelerate the forward pass of tensorized adapters by over 20%.
@@ 3.2 The AdaZeta Framework
MeZO(LoRA) 63.9 91.3 43.0 59.0 64.0 69.7 87.4
AdaZeta 64.3 91.5 49.6 60.7 68.1 68.7 87.5
@@ 3.3 Theoretical Analysis
In this subsection, we give the theoretical analysis for the AdaZeta framework. Our theoretical analysis highlights why the tensorized adapter and adaptive query schedule can significantly help to improve the ZO convergence rate. Unlike the theoretical analysis in the MeZO paper, which focuses on the ”effective rank” for the Hessian of loss, we focus on the dimension of the optimized models $d$ (number of trainable parameters) instead. As the trainable parameters with PEFT adapters are much smaller than the model size, the theoretical analysis based on the exact dimension of the optimization problem can better help us to explore the behavior of different PEFT methods.
@@ 4.1 Medium-size Roberta-Large Models
AdaZeta Shows Higher Accuracy than Other ZO Fine-Tuning Methods. According to our observations in Table 1, AdaZeta outperforms other ZO fine-tuning approaches in term of evaluation accuracy. Compared with MeZO-LoRA, which also involves PEFT adapters, AdaZeta outperforms in 6 out of 7 tests under both 16 and 64 batch size (BS) settings. This advantage shows the effectiveness of improving ZO estimation accuracy by further reducing the number of trainable parameters with the tensorized adapter. This is supported by the dimension-related convergence rate proved in Section 3.3.
@@ 4.2 Large-scale Llama-2 Models
To hilight the challenge of our experiments, we adopt a low-data resource approach using datasets from SuperGLUE Wang et al. (2019) and generative tasks such as SQuAD Rajpurkar et al. (2016) and DROP Dua et al. (2019). Our experimental protocol follows the prompted-based fine-tuning strategy outlined in the MeZO paper Malladi et al. (2023). The quantitive results are summarized in Table 2 and the training curves have been shown in Fig. 1. The following conclusions are drawn:
AdaZeta Method Demonstrates Superior Performance Over Traditional ZO Fine-Tuning. The AdaZeta framework delivers exceptional accuracy results across a variety of tasks, outperforming all ZO baseline methods such as MeZO and MeZO-LoRA in 8 out of 10 tasks. Compared with traditional inference-only methods like ICL and Zero-shot, AdaZeta significantly surpasses them with respect to test accuracy. Moreover, the AdaZeta method even outperforms the FO AdamW methods over several tasks like RTE, CB, and COPA, which require 8$\times$ more GPU memory.
AdaZeta Method Effectively Addresses Divergence Issues in ZO Fine-Tuning. We can observe from the table that the MeZO and MeZO-LoRA methods achieve unsatisfied results in some tasks like SST2, RTE, and BoolQ compared with our proposed method, which is led by the convergence issue. Also, we have shown that the AdaZeta method achieves lower evaluation loss much faster than the MeZO-LoRA and Sparse-MeZO methods across all tasks in Fig. 1. For example, the MeZO-LoRA method requires nearly 6K steps to achieve a loss of 0.4, whereas the AdaZeta method achieves the same degree of loss minimization in less than 1K steps, which represents a 6$\times$ speed-up with the same 1e-4 learning rate. Traditional ways to solve such divergence issues through increasing the batch size are hard to follow in the large-scale LLMs fine-tuning tasks. In contrast, the adaptive query schedule in the AdaZeta framework successfully mitigates this issue without increasing the training memory, thereby improving training outcomes.
[…]
@@ 4.3 Memory Training Time Efficiency
According to Fig. 3 (refer to Appendix B.1 for numberial results), the AdaZeta method requires only 14GB of memory to fine-tune the SST2 tasks on the Llama-2-7B model, which achieves over 8$\times$ Memory Reduction Relative to the FT Method. Also, compared with other ZO fine-tuning methods like MeZO, MeZO-LoRA, and Sparse-MeZO, the AdaZeta method utilizes similar or even less memory to achieve variance reduction. Traditional ways to reduce the ZO gradient estimation noise like increasing the batch size, consume significantly more memory than the AdaZeta method as shown in Fig. 3.
[…]
@@ 5 Conclusion
In this paper, we propose an adaptive zeroth-order fine-tuning framework with tensor-train decomposition, named AdaZeta. Compared with previous ZO fine-tuning works, the AdaZeta method achieves significantly better fine-tuning results across various tasks and models. Theoretical analysis has confirmed that our proposed methods enjoy better convergence, which is consistency with our experimental results on both Roberta-Large and Llama-2 models across various fine-tuning tasks.
@@
The primary limitation of this work is related to accelerating the proposed method. Currently, multiple queries at each training step are executed sequentially in a for-loop, which restricts further speed enhancements. This process can potentially be optimized by implementing parallel or distributed optimization techniques on GPUs, allowing for the simultaneous execution of multiple queries, as these queries are independent of each other with different random seeds. Also, Also, other gradient-free methods can be considered for this problem Hinton (2022); Mei et al. (2023) in different applications Zhang et al. (2024b); Cheng et al. (2023a)
Additionally, while we have introduced a fast-forward tensorized adapter in this study, further optimization is possible in the contraction process of these adapters during the forward pass. Advances in CUDA programming could significantly enhance the performance of our method. However, our current expertise in the domain of GPU optimization is limited, suggesting a potential area for collaboration or future research.
@@
This paper provides a cost-effective solution that operates with a minimal memory footprint. Event though we need to fine-tune large-scale models, the proposed method can alleviate the burden on data centers and reduce $CO_{2}$ emissions. However, we acknowledge that prolonged training times, especially with multiple GPUs, can pose environmental challenges. Consequently, our ongoing research endeavors are focused on developing more efficient training methods and preserving computational power with ecological considerations in mind.
@@ A.1 Dataset Setup
Table 4: Metrics that we use to evaluate the benchmark for Roberta-Large Model.
@@ B.1 Additional Momeory Comparison results
Table 9: Quantitive results for the memory profiling over SST2 and MultiRC tasks.
@@ Appendix C Proof of Theorem 1
where (a) can be proved with the use of the Lipschitz smoothness gradient denied in Assumption A1 that gives $x$ and $y$, we have $\|\nabla\ell(x)-\nabla\ell(y)\|\leq L\|x-y\|$. Additionaly, by the mean value theorem for vector-valued functions, there exists for any point $c$ on the line segment between $x$ and $y$ such that:
@@
Divide both side with $\eta$ and use the trick to introduce some randomly chosen $\bm{w}_{T}$ from the histroy with probability $P(T=k)=\frac{1}{K}$, we finish the proof as:

### v3

@@ Abstract
Fine-tuning large language models (LLMs) has achieved remarkable performance across various natural language processing tasks, yet it demands more and more memory as model sizes keep growing. To address this issue, the recently proposed Memory-efficient Zeroth-order (MeZO) methods attempt to fine-tune LLMs using only forward passes, thereby avoiding the need for a backpropagation graph. However, significant performance drops and a high risk of divergence have limited their widespread adoption. In this paper, we propose the Adaptive Zeroth-order Tensor-Train Adaption (AdaZeta) framework, specifically designed to improve the performance and convergence of the ZO methods. To enhance dimension-dependent ZO estimation accuracy, we introduce a fast-forward, low-parameter tensorized adapter. To tackle the frequently observed divergence issue in large-scale ZO fine-tuning tasks, we propose an adaptive query number schedule that guarantees convergence. Detailed theoretical analysis and extensive experimental results on Roberta-Large and Llama-2-7B models substantiate the efficacy of our AdaZeta framework in terms of accuracy, memory efficiency, and convergence speed. [^1] Code available on GitHub https://github.com/yifanycc/AdaZeta. [^2] Accepted by EMNLP 2024
@@
Yifan Yang1  Kai Zhen2  Ershad Banijamali2  Athanasios Mouchtaris2  Zheng Zhang1 1University of California, Santa Barbara 2Amazon AGI yifanyang@cs.ucsb.edu {kaizhen, ebanijam, mouchta}@amazon.com zhengzhang@ece.ucsb.edu
@@ 1 Introduction
Fine-tuning large language models (LLMs) has demonstrated outstanding performance in addressing numerous natural language processing applications, such as natural language understanding Kenton and Toutanova (2019), question-answering Xu et al. ; Cheng et al. (2023), and summarization Zhang et al. (2024). However, as the size of LLMs increases, the training process consumes progressively more GPU memory. In recent years, approaches such as quantization Tian et al. (2023); Dettmers et al. (2024) and parameter-efficient fine-tuning (PEFT) Hu et al. (2021) have been proposed to reduce memory costs during training by storing data with lower bit-depth or updating only a portion of the parameters. Despite these strategies effectively reducing memory costs, overall memory usage remains high due to the continuous reliance on a backpropagation graph.
@@ 3.2 The AdaZeta Framework
MeZO-LoRA 63.9 91.3 43.0 59.0 64.0 69.7 87.4
AdaZeta 64.3 91.5 49.6 60.7 68.1 68.7 86.5
@@ 3.3 Theoretical Analysis
In this subsection, we give the theoretical analysis for the AdaZeta framework. Our theoretical analysis highlights why the tensorized adapter and adaptive query schedule can significantly help to improve the ZO convergence rate. Unlike the theoretical analysis in the MeZO paper, which focuses on the ”effective rank” for the Hessian of loss, we focus on the dimension of the optimized models $d$ (number of trainable parameters) instead. As the trainable parameters with PEFT adapters are much smaller than the model size, the theoretical analysis based on the exact dimension of the optimization problem can better help us explore the behavior of different PEFT methods.
@@ 4.1 Medium-size Roberta-Large Models
AdaZeta Shows Higher Accuracy than Other ZO Fine-Tuning Methods. According to our observations in Table 1, AdaZeta outperforms other ZO fine-tuning approaches in terms of evaluation accuracy. Compared with MeZO-LoRA, which also involves PEFT adapters, AdaZeta outperforms in 5 out of 7 tests under both 16 and 64 batch size (BS) settings. This advantage shows the effectiveness of improving ZO estimation accuracy by further reducing the number of trainable parameters with the tensorized adapter. This is supported by the dimension-related convergence rate proved in Section 3.3.
@@ 4.2 Large-scale Llama-2 Models
To highlight the challenge of our experiments, we adopt a low-data resource approach using datasets from SuperGLUE Wang et al. (2019) and generative tasks such as SQuAD Rajpurkar et al. (2016) and DROP Dua et al. (2019). Our experimental protocol follows the prompted-based fine-tuning strategy outlined in the MeZO paper Malladi et al. (2023). The quantitative results are summarized in Table 2 and the training curves have been shown in Fig. 1. Note that it is reasonable to observe some large accuracy gap between different methods under different tasks, which has also been observed in previous MeZO and PEFT papers Malladi et al. (2023); Hu et al. (2023). The following conclusions are drawn:
AdaZeta Method Demonstrates Superior Performance Over Traditional ZO Fine-Tuning. The AdaZeta framework delivers exceptional accuracy results across a variety of tasks, outperforming all ZO baseline methods such as MeZO and MeZO-LoRA in 8 out of 10 tasks. Compared with traditional inference-only methods like ICL and Zero-shot, AdaZeta significantly surpasses them with respect to test accuracy. Moreover, the AdaZeta method even outperforms the FO-AdamW methods over several tasks like RTE, CB, and COPA, which require 8$\times$ more GPU memory.
AdaZeta Method Effectively Addresses Divergence Issues in ZO Fine-Tuning. We can observe from the table that the MeZO and MeZO-LoRA methods achieve unsatisfied results in some tasks like SST2, RTE, and BoolQ compared with our proposed method, which is led by the convergence issue. Also, we have shown that the AdaZeta method achieves lower evaluation loss much faster than the MeZO-LoRA and Sparse-MeZO methods across all tasks in Fig. 1. For example, the MeZO-LoRA method requires nearly 6K steps to achieve a loss of 0.4, whereas the AdaZeta method achieves the same degree of loss minimization in less than 1K steps, which represents a 6$\times$ speed-up with the same 1e-4 learning rate. Traditional ways to solve such divergence issues through increasing the batch size are hard to follow in the large-scale LLMs fine-tuning tasks. In contrast, the adaptive query schedule in the AdaZeta framework successfully mitigates this issue without increasing the training memory, thereby improving training outcomes. Additionally, we observed that combining LoRA with the adaptive query schedule significantly improves performance in certain tasks. Future work could also explore incorporating the adaptive query schedule into the MeZO-LoRA method to further enhance stability.
Table 3: Required GPU hours (GPU numbers $\times$ Training hours) to achieve each evaluation loss for different ZO fine-tuning methods on Llama-2-7B model.
@@ 4.3 Memory Training Time Efficiency
According to Fig. 3 (refer to Appendix B.1 for numerical results), the AdaZeta method requires only 14GB of memory to fine-tune the SST2 tasks on the Llama-2-7B model, which achieves over 8$\times$ Memory Reduction Relative to the FT Method. Also, compared with other ZO fine-tuning methods like MeZO, MeZO-LoRA, and Sparse-MeZO, the AdaZeta method utilizes similar or even less memory to achieve variance reduction. Traditional ways to reduce the ZO gradient estimation noise like increasing the batch size, consume significantly more memory than the AdaZeta method as shown in Fig. 3.
In Table 3, we measure the total GPU hours required to achieve a certain threshold of training loss across four tasks (SST2, WIC, CB, MultiRC). For the applicability of the experiments, we established an evaluation loss threshold that all methods could achieve. According to the results, it is evident that the AdaZeta method converges on-par or faster than other ZO fine-tuning methods with even better results than the MeZO-LoRA and Sparse-MeZO methods under the large-batch size case. Note that we did not utilize the gradient accumulation technique for the 64 batch size case, which may significantly increase the training time.
Table 4: Compare with first-order LoRA method under low ranks and batch sizes.
Setup LoRA/r=1/BS=1 LoRA/r=1/BS=8 LoRA/r=8/BS=8 AdaZeta/r=8/BS=1 MeZO-LoRA/r=8/BS=16 AdaZeta/r=8/BS=16
Memory (GB) 35.60 96.65 96.72 14.05 23.02 23.01
4.4 Further Comparison with LoRA
In this section, we further compare our AdaZeta method with the first-order LoRA method in terms of training memory usage across different ranks and batch sizes. The results for the CB task are presented in Table 4. We make the following observations under two scenarios:
Reducing the LoRA Rank: Reducing the LoRA rank (even down to 1) has minimal impact on training memory in the first-order setting. The reason is that the backpropagation graph—which contains intermediate gradient information—still needs to be retained, spanning almost the entire model in the vanilla LoRA approach.
Reducing the Batch Size: Reducing the batch size is a more effective way to reduce the training memory for both FO and ZO cases. With the existence of a backpropagation graph, it is reasonable to observe a larger reduction of training memory of the FO method than ZO when reducing the number of batch sizes. However, we can observe that even when comparing our method with the LoRA method using a batch size of 1, our method is still 2.5$\times$ more memory-efficient. Additionally, even comparing AdaZeta/r=8/BS=16 with LoRA/r=1/BS=1, we still achieve nearly a 50% reduction in memory usage. However, we would like to remark that the batch size of 1 setup is rarely used in practice due to the following reasons:
  • First, reducing the batch size will dramatically increase the training time of the LoRA method.
  • Second, such a small batch size leads to large stochastic noise during the fine-tuning process, which further harms the training performance. Hu et al. (2023)
@@ 5 Conclusion
In this paper, we propose an adaptive zeroth-order fine-tuning framework with tensor-train decomposition, named AdaZeta. Compared with previous ZO fine-tuning works, the AdaZeta method achieves significantly better fine-tuning results across various tasks and models. Theoretical analysis has confirmed that our proposed methods enjoy better convergence, which is consistent with our experimental results on both Roberta-Large and Llama-2 models across various fine-tuning tasks.
Future work could explore improving the efficiency of the AdaZeta method by implementing distributed optimization across multiple GPUs for handling multiple queries concurrently at each step. Additionally, applying the adaptive query schedule to other PEFT methods may yield significantly better performance compared to the original MeZO algorithm.
Acknowledgements
This project was supported by Amazon. We extend our gratitude to Siegfried Kunzmann, Jiajun Zhou, Clement Chung, Samridhi Choudhary, Hieu Nguyen and the many other colleagues at Amazon AGI and UCSB who engaged in discussions that shaped this work.
This research also utilized resources from the National Energy Research Scientific Computing Center (NERSC), a U.S. Department of Energy Office of Science User Facility, supported under Contract No. DE-AC02-05CH11231 through NERSC award ASCR-ERCAP0030039.
@@ Acknowledgements
The primary limitation of this work is related to accelerating the proposed method. Currently, multiple queries at each training step are executed sequentially in a for-loop, which restricts further speed enhancements. This process can potentially be optimized by implementing parallel or distributed optimization techniques on GPUs, allowing for the simultaneous execution of multiple queries, as these queries are independent of each other with different random seeds.
@@
This paper provides a cost-effective solution that operates with a minimal memory footprint. Even though we need to fine-tune large-scale models, the proposed method can alleviate the burden on data centers and reduce $CO_{2}$ emissions. However, we acknowledge that prolonged training times, especially with multiple GPUs, can pose environmental challenges. Consequently, our ongoing research endeavors are focused on developing more efficient training methods and preserving computational power with ecological considerations in mind.
@@ A.1 Dataset Setup
Table 5: Metrics that we use to evaluate the benchmark for the Roberta-Large Model.
@@ B.1 Additional Momeory Comparison results
Table 10: Quantitative results for the memory profiling over SST2 and MultiRC tasks.
@@ Appendix C Proof of Theorem 1
where (a) can be proved with the use of the Lipschitz smoothness gradient denied in Assumption A1 that gives $x$ and $y$, we have $\|\nabla\ell(x)-\nabla\ell(y)\|\leq L\|x-y\|$. Additionally, by the mean value theorem for vector-valued functions, there exists for any point $c$ on the line segment between $x$ and $y$ such that:
@@
Divide both side with $\eta$ and use the trick to introduce some randomly chosen $\bm{w}_{T}$ from the history with probability $P(T=k)=\frac{1}{K}$, we finish the proof as: